can you help me answer this question
2(5x-1)=4-3x/2

Found 2 solutions by atif.muhammad, jen0921:
Answer by atif.muhammad(135)

Are you sure this is a quadratic equation?
x + 3 + 2x = x + 3
In order to call it a quadratic equation it must have an "x^2"

The solution to the equation you've given me is:
x + 3 + 2x = x+ 3
3x + 3 = x + 3
2x = 0
x= 0

Answer by jen0921(1)